OpenStudy (anonymous):

Craig started running track after school to get ready for basketball tryouts. The first day, he ran the quarter-mile track six times in 15 minutes. What was his rate of speed? I think its 2 or 3 miles per hour

OpenStudy (anonymous):

@d3Xter

OpenStudy (anonymous):

i seriously dont know. i m very weak at this subject sory i cant help u

OpenStudy (anonymous):

thats okay

OpenStudy (jack1):

speed = distance / time

OpenStudy (jack1):

so distance = he ran the quarter-mile track six times quarter mile = 1/4 miles = 0.25 miles 6 times so 6 x 0.25 miles = 1.5 miles

OpenStudy (jack1):

time = 15 minutes = 1/4 hours = 0.25 hours so speed = distance / time = 1.5 miles / 0.25 hours = 6 miles / hour

OpenStudy (anonymous):

okay thanks I think I understand know

OpenStudy (jack1):

hall good dude, always best to first convert to the units u want everything in so in this case mph